OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of -6 and a y-intercept of 9.

OpenStudy (anonymous):

y = mx + c where m is slope and c is y intercept.. Plug in the values..

OpenStudy (saifoo.khan):

y = mx + b Where m is the slope and b is the y-intercept. y = -6x + 9

OpenStudy (anonymous):

y=mx+c

OpenStudy (anonymous):

So: y = -6x + 9

OpenStudy (anonymous):

@waterineyes i agree

OpenStudy (anonymous):

And I agree with you too @laxplayer